The amount of money that a salesperson makes varies as the total amount of sales made.If the sales person receives $7,500 for selling a house for $120,000 how much will he make if he sells a house for $ 150,000?
Found 2 solutions by checkley71, ikdeep:
Answer by checkley71(8403)

7500/120000=x/15000 or120x=7500*15000 or 120000x=112500000 or x=9375

Answer by ikdeep(226)

You can also find solution to this problem through percentage system...and is goes like this
now Let sales person percentage commission on selling a house = x%
and acoording to question ,,i.e. sales person receives $7,500 for selling a house for $120,000 ...we have get the equation..
(7,500/120,000) * 100 = x
on solving it we get
x = 6.25%
Now to calculate his commission on selling house of $150,000 = 150,000 * 6.25%
= $9,375